OpenStudy (anonymous):

Identify the standard form of the equation 4y^2-338x+32y=-169x^2+443

OpenStudy (campbell_st):

well it looks like a hyperbola... collect like terms \[4y^2 + 32y + 169x^2 -338x = 443\]

OpenStudy (anonymous):

I did that but i dont know what to do next

OpenStudy (campbell_st):

ok... so do this so complete the square for both x and y \[4y^2 + 32y + 16 + 169x^2 - 338x + 169 = 443 - 16 - 169\] which becomes \[(2y + 4)^2 + (13x - 13)^2 = 258\] hope it helps